61. (14.8) Find the extreme values for the function 8y - 4:subject to the constraints y2 + x2 - 1 = 0 and 2x - z-y-2 = 0.
Answer #1

let f(x,y)= 84 - 4x 9(8,9,2) = 27x²1=0 h (2,9,2) = 2x-2-4-2=0 all want to ind extreme value we of f(x, y, 2) of thash Subjectd 외 85 | TE A = 2.55, hon x- 13-2 2 .) - - , Z : -409 | 50( 45) 1. 15. k u- -2 J5 ) ( 55- -4플용 여1훓1볼 (무) -() 48 = CG, 플무) Is